# Influence of Camera-LiDAR Configuration on 3D Object Detection for Autonomous Driving

### v1

@@ Abstract
Cameras and LiDARs are both important sensors for autonomous driving, playing critical roles for 3D object detection. Camera-LiDAR Fusion has been a prevalent solution for robust and accurate autonomous driving perception. In contrast to the vast majority of existing arts that focus on how to improve the performance of 3D target detection through cross-modal schemes, deep learning algorithms, and training tricks, we devote attention to the impact of sensor configurations on the performance of learning-based methods. To achieve this, we propose a unified information-theoretic surrogate metric for camera and LiDAR evaluation based on the proposed sensor perception model. We also design an accelerated high-quality framework for data acquisition, model training, and performance evaluation that functions with the CARLA simulator. To show the correlation between detection performance and our surrogate metrics, We conduct experiments using several camera-LiDAR placements and parameters inspired by self-driving companies and research institutions. Extensive experimental results of representative algorithms on NuScenes dataset validate the effectiveness of our surrogate metric, demonstrating that sensor configurations significantly impact point-cloud-image fusion based detection models, which contribute up to 30% discrepancy in terms of average precision.
@@ I Introduction
Fast evaluation of 3D detection performance under different camera and LiDAR configurations in the real world is quite challenging due to the laboriousness of data acquisition, model training, and performance testing [12]. Besides, efficiently comparing different sensor configurations for better 3D perception remains an open and critical question under a general trend of using more multi-modal sensors in autonomous driving [19]. To this point, this paper investigates the impact of camera-LiDAR configuration on 3D object detection performance, and proposes a novel and unified framework for accelerating the evaluation of different camera-LiDAR configurations. The main contributions of this paper are summarized as follows:
  • We establish a new systematic framework to efficiently evaluate the 3D detection performance of different camera-LiDAR configurations without the effort-costly loop of data collection, model training and evaluation.
@@
  • Extensive experimental results in CARLA validate the correlation between our surrogate metric and the detection performance of several camera-LiDAR fusion algorithms. The code is available on https://github.com/safeai-lab/lidar-camera-placement.
@@ II-A Multi-modal Sensor Configurations
Perception is an important sub-module of the autonomous driving system [20], which directly affects the decision-making and behavior of vehicles. With this respect, LiDARs and cameras are widely used for the perception of autonomous vehicles due to their capability to obtain rich information from the environment in real-time setting [21, 22, 23]. However, the performance of cameras and LiDARs is sensitive to the physical installation errors or motion perturbations [17, 24, 25]. There exist some works exploring configurations of cameras and LiDARs separately. Rahimian et al. [26] introduce a dynamic occlusion optimization method to estimate the configuration of the single camera.Puligandla et al. [15] optimize multi-camera placements for vehicle surround view in continuous domain using gradient-free black-box optimization. For LiDARs, Liu et al. [27] develop a non-detectable-space-based surrogate function and realize the optimum with Artificial Bee Colony algorithm. Instead of optimizing the surrounding view of sensors on the vehicle and to minimize undetectable space, recent works [12] quantitatively investigates the interaction between LiDAR placement and detection performance and proposes metrics to evaluate LiDAR perception. However, quantitative evaluation methods for both cameras and LiDARs associated with sensor fusion based 3D detection algorithms remain to be explored although [14] initiates the study of multiple sensors empirically. To this end, we present a novel unified metric for quantitatively studying both camera and LiDAR configurations in Section III.
@@ II-B Camera-LiDAR Detection Algorithms
3D object detection is an important task in autonomous driving. Recently, camera-LiDAR fusion has attracted increasing attention to boost more robust and accurate detection performance. Current camera-LiDAR fusion work can be classified into three categories: result-level, proposal-level, and point-level. The result-level methods [28, 29, 30, 31] use 2D object detectors to prepare 3D detection proposals, with difficulty encountered when objects overlap in 2D planes. The proposal-level methods [32, 33] generate 3D proposals directly and perform fusion at the region-of-interest (ROI) level through ROIPool [34], but the granularity is poor. Most recent works use the point-level method and achieve promising detection performance. Point-level fusion methods established the association between 3D points and image pixels through calibration matrices and decorate LiDAR points input with image segmentation features, including input-level decoration [35, 9, 36, 37, 38, 39] and feature-level decoration. Moreover, some approaches [40, 41, 42] perform the decoration process on the bird’s eye view (BEV) plane. The majority of the detection methods mentioned above are well-designed and evaluated on high-quality point cloud datasets, but they do not consider the influence of the placements and parameters of the camera and LiDAR sensing system. We adopt several representative latest released methods to evaluate the influence of camera-LiDAR configurations on detection performance in Section IV.
@@ III-A Problem Formulation
In our work, ROI refers to the limited 3D cuboid area $[L,W,H]$, where the sensor detects the target object. We divide the 3D ROI cuboid area $\Omega$ into voxels of the same resolution $\delta$,
@@ III-B Modeling Camera-LiDAR Perception
In our camera perception model, the number of rays does not exactly equal to the original pixel resolution $w_{0}\times h_{0}$ The images obtained by the camera are dense and with high-resolution, but learning-based methods do not directly take the original images as inputs without down-sampling. So the original images are resized to a lower resolution $u_{0}\times v_{0}$ to model the process of the camera ray casting. We first map the pixels on the resized image to the pixels on the original image as follows,
@@ III-C Multi-sensor Probabilistic Occupancy Grid
where $\omega_{i}\sim p_{\Omega}$. Given a dataset with ground true of samples $Y_{T}=\{y_{1},y_{2},...,y_{T}\}$, $T$ denotes the total frames of ground truth samples. Each frame $y_{t}\in\{y_{1},...,y_{T}\}$ of the given samples contains $S^{(t)}$ ground-truth bounding boxes. We denote the bounding boxes of target objects as $\{b_{1}^{(t)},b_{2}^{(t)},...,b_{S^{(t)}}^{(t)}\}$. We use $\omega_{i}\in y_{t}$ to denote the case where the voxel $\omega_{i}$ is contained by the ground truth bounding box of any target object in the frame $y_{t}$. Then, for each voxel $\omega_{i}$ in the ROI, we traverse all the $T$ frames of samples to obtain the probability that the voxel $\omega_{i}$ is occupied by the target objects in any given frame from $Y_{T}$ as,
$\hat{p}(\omega_{i})=\frac{\sum_{t=1}^{T}\mathds{1}(\omega_{i}\in y_{t})}{T},~{}\omega_{i}\in y_{t}:=\exists\;b^{(t)}\in y_{t},s.t.\;\omega_{i}\in b^{(t)}$
where $\mathds{1}(\cdot)$ is an indicator function and $N$ is the number of voxels in ROI. The POG can then be estimated by the joint probability of all occupied voxels in ROI. Since the presence of an object in one voxel does not imply presence in other voxels among all the frames in ROI, we could treat these voxels as independent and identically distributed random variables and calculate the joint distribution over all voxels in the set $\Omega$ as,
@@
where $N$ is the total number of non-zero voxels in ROI. Note that notations of $\hat{p}$ with $\hat{}$ are the estimated distribution from observed samples, while notations of $p$ without $hat$ are the unknown non-random statistics to be estimated.
@@
The conditional Probabilistic Occupancy Grid reflects the conditional joint distribution of voxels that intersect with the perceptual area of the sensor in the ROI. From the perspective of density estimation to find POG, the true POG and conditional POG given configuration $C_{0}$ can be estimated as
$p_{POG}=\hat{p}_{POG},p_{POG|C_{0}}=\hat{p}_{POG|C_{0}}$ (11)
@@ III-D Unified Surrogate Metric for Multi-sensor Fusion
Perception performance of a given configuration is maximized when the uncertainty of the voxel distribution to the sensor is minimized. We denote the information gain $IG$ using the mutual information (MI) between the full entropy and the conditional entropy, which represents the decrease of the uncertainty of the voxel distribution given a specific sensor configuration.
@@
$S_{MS}=S_{MIG|C^{c}(\theta^{c},\psi^{c})}+\lambda S_{MIG|C^{L}(\theta^{L},\psi^{L})}$ (16)
where $\lambda$ is a hyper-parameter weighting the difference in perceptual capbility between LiDAR and the camera, to fairly show the influence of cameras and LiDARs on 3D detection.
@@ IV Experiments
In this section, we design a framework for automatic camera-LiDAR data collection and evaluation based on the CARLA simulator [45] to avoid time-consuming, low-efficiency and high-cost real-world experiments to validate our method. To ensure the fairness of the experimental comparison, all images and point clouds collected in CARLA are with fixed scenarios, including the route of the data-collection ego vehicle, driving scenarios, traffic flow, etc. We conducted comprehensive experiments to show two key points: the impact of camera and LiDAR configurations on 3D object detection performance, and the correlation between our proposed unified surrogate metric and learning-based perception performance.
@@ IV-A Experimental Setup
Camera-LiDAR fusion detection algorithms and metrics. In order to fairly demonstrate the performance under different camera-LiDAR configurations on 3D object detection, we adopt the three representative open-source camera-LiDAR fusion detection method. Specifically, we use two two-stage detection methods, Transfusion [7] and Bevfusion [8], and one single-stage detection method, Pointaugmenting [9]. We follow the default training pipeline and hyperparameters of these methods. For two-stage methods, We first train the LiDAR-only detection model, then fine-tune the fusion methods with pretrained image detection model. For the one-stage method, we train two backbones, VoxelNet [47], and Pointpillars [48], respectively. We adopt the mean Average Precision (mAP) metric from NuScenes [46], which defines thresholds by considering the 2D distance on the ground rather than Intersection over Union (IoU) based ones [21]. For mAP, The 2D distances are at 4 different thresholds: 0.5, 1.0, 2.0, and 4.0, respectively.
Different sensor configurations. The sensor configurations in this work adopt 4 LiDARs and 6 cameras following the NuScenes dataset [49]. Four different LiDAR configurations inspired by famous autonomous driving companies [12] and two camera configurations are used to explore the influence on object detection performance. For LiDAR configurations, the beams are be equally distributed in the vertical FOV [-25.0, 25.0] degrees, as shown in Fig. 1. Center placement is achieved by vertically stacking four LiDARs together at the roof center (1a). Pyramid placement (Fig. 1b) includes 1 front LiDAR and 3 back ones, with a higher one in the middle. Line has (1c) 4 LiDARs placed in a horizontal line symmetrically. Trapezoid installs 4 LiDARs in the parallel front and back (Fig. 1d). Each camera has the same resolution 1600×900. Wide configuration is the same as the setup of NuScenes with 5 general cameras (FOV = 70°) and a wide-angle camera (FOV = 110°) on the roof of the ego car (Fig. 1e). As contrast to the wide one, we design another Narrow configuration by changing the FOV to 55 degrees for all cameras in Fig. 1f.
@@ IV-B Experiment Results and Analysis
Influence of sensor configurations on 3D object detection. In Fig. 5 and Table II, we show the 3D target detection performance with different representative algorithms using different camera-LiDAR configurations. The configurations of sensors drastically influence the detection performance, with a maximum fluctuation of 30%. Specifically, for the effects of the camera and LiDAR configurations respectively, it is not difficult to find that for the same LiDAR configuration, the Wide camera configurations significantly outperform the Narrow camera configurations; and for the same camera configuration, Trapezoid has the best detection performance for the majority of algorithms and objects. Under the Line LiDAR configuration, Pointaugmenting is the best for Bicycle while Bevfusion is the best for the Pedestrian, due to the point cloud detection characteristics of specific algorithms and the influence of POG for different objects.
@@
Potential application analysis. Our unified surrogate metric greatly accelerates the development, optimization, and evaluation of multi-sensor configurations for self-driving cars compared to the commonly-seen R&D process involving installing sensors, collecting data, training models, and evaluating performance. Our approach can efficiently evaluate different camera-LiDAR configurations by simply calculating the surrogate metric from the bounding boxes. Besides, the optimal solution of camera-LiDAR configurations for specific scenarios can be found based on our metrics.

### v2

@@ Abstract
Cameras and LiDARs are both important sensors for autonomous driving, playing critical roles in 3D object detection. Camera-LiDAR Fusion has been a prevalent solution for robust and accurate driving perception. In contrast to the vast majority of existing arts that focus on how to improve the performance of 3D target detection through cross-modal schemes, deep learning algorithms, and training tricks, we devote attention to the impact of sensor configurations on the performance of learning-based methods. To achieve this, we propose a unified information-theoretic surrogate metric for camera and LiDAR evaluation based on the proposed sensor perception model. We also design an accelerated high-quality framework for data acquisition, model training, and performance evaluation that functions with the CARLA simulator. To show the correlation between detection performance and our surrogate metrics, We conduct experiments using several camera-LiDAR placements and parameters inspired by self-driving companies and research institutions. Extensive experimental results of representative algorithms on nuScenes dataset validate the effectiveness of our surrogate metric, demonstrating that sensor configurations significantly impact point-cloud-image fusion based detection models, which contribute up to 30% discrepancy in terms of the average precision.
@@ I Introduction
Fast evaluation of 3D detection performance under different camera and LiDAR configurations in the real world is quite challenging due to the laboriousness of data acquisition, model training, and performance testing [12]. Besides, efficiently comparing different sensor configurations for better 3D perception remains an open and critical question under a general trend of using more multi-modal sensors in autonomous driving [19]. To this point, this paper investigates the impact of camera-LiDAR configuration on 3D object detection performance and proposes a novel and unified framework for accelerating the evaluation of different camera-LiDAR configurations. The main contributions of this paper are summarized as follows:
  • We establish a new systematic framework to efficiently evaluate the 3D detection performance of different camera-LiDAR configurations without the effort-costly loop of data collection, model training, and evaluation.
@@
  • Experimental results in CARLA validate the correlation between our unified surrogate metric and the performance of several camera-LiDAR algorithms. The code is available on https://github.com/ywyeli/lidar-camera-placement.
@@ II-A Multi-modal Sensor Configurations
Perception is an important sub-module of the autonomous driving system [20], which directly affects the decision-making and behavior of vehicles. With this respect, LiDARs and cameras are widely used for the perception of autonomous vehicles due to their capability to obtain rich information from the environment in real-time settings [21, 22, 23]. However, the performance of cameras and LiDARs is sensitive to physical installation errors or motion perturbations [17, 24, 25]. There exist some works exploring configurations of cameras and LiDARs separately. Rahimian et al. [26] introduce a dynamic occlusion optimization method to estimate the configuration of the single camera. Puligandla et al. [15] optimize multi-camera placements for vehicle surround view in continuous domain using gradient-free black-box optimization. For LiDARs, Liu et al. [27] develop a non-detectable-space-based surrogate function and realize the optimum with the Artificial Bee Colony algorithm. Instead of optimizing the surrounding view of sensors on the vehicle to minimize undetectable space, recent works [12] quantitatively investigate the interaction between LiDAR placement and detection performance and propose metrics to evaluate LiDAR perception. However, quantitative evaluation methods for both cameras and LiDARs associated with sensor fusion based 3D detection algorithms remain to be explored although [14] initiates the study of multiple sensors empirically. To this end, we present a novel unified metric for quantitatively studying both camera and LiDAR configurations in Section III.
@@ II-B Camera-LiDAR Detection Algorithms
3D object detection is an important task in autonomous driving. Recently, camera-LiDAR fusion has attracted increasing attention to boost more robust and accurate detection performance. Current camera-LiDAR fusion work can be classified into three categories: result-level, proposal-level, and point-level. The result-level methods [28, 29, 30, 31] use 2D object detectors to prepare 3D detection proposals, with difficulty encountered when objects overlap in 2D planes. The proposal-level methods [32, 33] generate 3D proposals directly and perform fusion at the region-of-interest (ROI) level through ROIPool [34], but the granularity is poor. Most recent works use the point-level method and achieve promising detection performance. Point-level fusion methods established the association between 3D points and image pixels through calibration matrices and decorate LiDAR points input with image segmentation features, including input-level decoration [35, 9, 36, 37, 38, 39] and feature-level decoration. Moreover, some approaches [40, 41, 42] perform the decoration process on the bird’s eye view (BEV) plane. The majority of the detection methods mentioned above are well-designed and evaluated on high-quality point cloud datasets, but they do not consider the placements of the camera-LiDAR sensing system. We adopt several representative latest methods to evaluate the influence of camera-LiDAR configurations on detection performance in Section IV.
@@ III-A Problem Formulation
In our work, ROI refers to the limited 3D cuboid area $[L,W,H]$, where the sensor detects the objects. We divide the ROI area $\Omega$ into voxels of the same resolution $\delta$,
@@ III-B Modeling Camera-LiDAR Perception
In our camera perception model, the number of rays does not exactly equal the original pixel resolution $w_{0}\times h_{0}$ The images obtained by the camera are dense and with high resolution, but learning-based methods do not directly take the original images as inputs without down-sampling. So the original images are resized to a lower resolution $u_{0}\times v_{0}$ to model the process of the camera ray casting. We first map the pixels on the resized image to the pixels on the original image as follows,
@@ III-C Multi-sensor Probabilistic Occupancy Grid
where $\omega_{i}\sim p_{\Omega}$ and $N$ is the number of voxels in ROI. Given a dataset with ground true of samples $Y_{T}=\{y_{1},y_{2},...,y_{T}\}$, $T$ denotes the total frames of ground truth samples. Each frame $y_{t}\in\{y_{1},...,y_{T}\}$ of the given samples contains $S^{(t)}$ ground-truth bounding boxes. We denote the bounding boxes of target objects as $\{b_{1}^{(t)},b_{2}^{(t)},...,b_{S^{(t)}}^{(t)}\}$. We use $\omega_{i}\in y_{t}$ to denote the case where the voxel $\omega_{i}$ is contained by the ground truth bounding box of any target object in the frame $y_{t}$. Then, for each voxel $\omega_{i}$ in the ROI, we traverse all the $T$ frames of samples to obtain the probability that the voxel $\omega_{i}$ is occupied by the target objects in any given frame from $Y_{T}$ as,
$\hat{p}(\omega_{i})=\frac{\sum_{t=1}^{T}\mathds{1}(\omega_{i}\in y_{t})}{T},$ (7)
$\omega_{i}\in y_{t}:=\exists\;b^{(t)}\in y_{t},\;s.t.\;\omega_{i}\in b^{(t)}$
where $\mathds{1}(\cdot)$ is an indicator function. The POG can then be estimated by the joint probability of all occupied voxels in ROI. Since the presence of an object in one voxel does not imply presence in other voxels among all the frames in ROI, we could treat these voxels as independent and identically distributed random variables and calculate the joint distribution over all non-zero voxels in the set $\Omega$ as,
@@
where $N$ is the total number of voxels in ROI. Note that notations of $\hat{p}$ with $\hat{}$ are the estimated distribution from observed samples, while notations of $p$ without $hat$ are the unknown non-random statistics to be estimated.
@@
The conditional POG reflects the conditional joint distribution of voxels that intersect with the perceptual area of the sensor in the ROI. From the perspective of density estimation to find POG, the true POG and conditional POG given configuration $C_{0}$ can be estimated as
$p_{POG}=\hat{p}_{POG},\\ \ p_{POG|C_{0}}=\hat{p}_{POG|C_{0}}$ (12)
@@ III-D Unified Surrogate Metric for Multi-sensor Fusion
Perception performance of a given configuration is maximized when the uncertainty of the voxel distribution to the sensor is minimized[12]. We denote the information gain $IG$ using the mutual information (MI) between the full entropy and the conditional entropy, which represents the decrease of the uncertainty of the voxel distribution given a specific sensor configuration.
@@
$S_{MS}=\lambda S_{MIG|C^{c}(\theta^{c},\psi^{c})}+S_{MIG|C^{L}(\theta^{L},\psi^{L})}$ (17)
where $\lambda$ is a hyper-parameter weighting the difference in perceptual capability between LiDAR and the camera, to fairly show the influence of cameras and LiDARs on 3D detection.
@@ IV Experiments
In this section, we design a framework for automatic camera-LiDAR data collection and evaluation based on the CARLA simulator [45] to avoid time-consuming, low-efficiency, and high-cost real-world experiments to validate our method. To ensure the fairness of the experimental comparison, all images and point clouds collected in CARLA are with fixed scenarios, including the route of the data-collection ego vehicle, driving scenarios, traffic flow, etc. We conducted comprehensive experiments to show two key points: the impact of camera and LiDAR configurations on 3D object detection performance, and the correlation between our proposed unified surrogate metric and learning-based perception performance.
@@ IV-A Experimental Setup
Camera-LiDAR fusion detection algorithms and metrics. To fairly demonstrate the performance under different camera-LiDAR configurations on 3D object detection, we adopt the three representative open-source camera-LiDAR fusion detection methods. Specifically, we use two two-stage detection methods, Transfusion [7] and Bevfusion [8], and one single-stage detection method, Pointaugmenting [9]. We follow the default training pipeline and hyperparameters of these methods. For two-stage methods, We first train the LiDAR-only detection model, then fine-tune the fusion methods with pretrained image detection models. For the one-stage method, we train two backbones, VoxelNet [47], and Pointpillars [48], respectively. We adopt the mean Average Precision (mAP) metric from nuScenes [46], which defines thresholds by considering the 2D distance on the ground rather than Intersection over Union (IoU) based ones [21].
Different sensor configurations. The sensor configurations in this work adopt 4 LiDARs and 6 cameras following nuScenes [49]. Four different LiDAR configurations inspired by famous autonomous driving companies [12] and two camera configurations are used to explore the influence on object detection performance. For LiDAR configurations, the beams are equally distributed in the vertical FOV [-25.0, 25.0] degrees, as shown in Fig. 1. Center placement is achieved by vertically stacking four LiDARs together at the roof center (1a). Pyramid placement (Fig. 1b) includes 1 front LiDAR and 3 back ones, with a higher one in the middle. Line has (1c) 4 LiDARs placed in a horizontal line symmetrically. Trapezoid installs 4 LiDARs in the parallel front and back (Fig. 1d). Each camera has the same resolution 1600×900. Wide configuration is the same as the setup of nuScenes with 5 general cameras (FOV = 70°) and a wide-angle camera (FOV = 110°) on the roof of the ego car (Fig. 1e). As contrast to the Wide configuration, we design another Narrow configuration by changing the FOV to 55 degrees for all cameras in Fig. 1f.
@@ IV-B Experiment Results and Analysis
Influence of sensor configurations on 3D object detection. In Fig. 5 and Table II, we show the 3D target detection performance with different representative algorithms using different camera-LiDAR configurations. The configurations of sensors drastically influence the detection performance, with a maximum fluctuation of 30%. Specifically, for the effects of the camera and LiDAR configurations respectively, it is not difficult to find that for the same LiDAR configuration, the Wide camera configurations significantly outperform the Narrow camera configurations for up to 10%; and for the same camera configuration, Trapezoid has the best detection performance for the majority of algorithms and objects. Under the Line LiDAR configuration, Pointaugmenting is the best for Bicycle while Bevfusion is the best for the Pedestrian, due to the point cloud detection characteristics of specific algorithms and the influence of POG for different objects.
@@
Potential application analysis. Our unified surrogate metric greatly accelerates the development, optimization, and evaluation of multi-sensor configurations for self-driving cars compared to the commonly seen R&D process involving installing sensors, collecting data, training models, and evaluating performance. Our approach can efficiently evaluate different camera-LiDAR configurations by simply calculating the surrogate metric from the bounding boxes. Besides, the optimal solution of camera-LiDAR configurations for specific scenarios can be found based on our metrics.
Future work. The unified surrogate metric could be extended for more algorithms and include additional sensors[52], such as solid-state LiDAR and radar. Experiments on real sensors and the latest placements from the autonomous driving industry [53, 54] are crucial for enhancing the effectiveness of our metric. The comparison of performance on data from simulation platforms and real-world data also represents an exciting frontier. Further exploration of optimizing the surrogate metric score to enhance the LiDAR placement would provide valuable insights.